# Get away with less: Need of source side data curation to build parallel corpus for low resource Machine Translation

### v1

@@ Abstract
Data curation is a critical yet under-researched step in the machine translation training paradigm. To train translation systems, data acquisition relies primarily on human translations and digital parallel sources or, to a limited degree, synthetic generation. But, for low-resource languages, human translation to generate sufficient data is prohibitively expensive. Therefore, it is crucial to develop a framework that screens source sentences to form efficient parallel text, ensuring optimal MT system performance in low-resource environments. We approach this by evaluating English-Hindi bi-text to determine effective sentence selection strategies for optimal MT system training. Our extensively tested framework, (Lexical And Linguistically Informed Text Analysis) LALITA, targets source sentence selection using lexical and linguistic features to curate parallel corpora. We find that by training mostly on complex sentences from both existing and synthetic datasets, our method significantly improves translation quality. We test this by simulating low-resource data availabilty with curated datasets of 50K to 800K English sentences and report improved performances on all data sizes. LALITA demonstrates remarkable efficiency, reducing data needs by more than half across multiple languages (Hindi, Odia, Nepali, Norwegian Nynorsk, and German). This approach not only reduces MT systems training cost by reducing training data requirement, but also showcases LALITA’s utility in data augmentation.
@@ 1 Introduction
To address this scarcity, foundational efforts have focused on creating and managing large-scale corpora from the ground up. For many low-resource languages, like Mizo (one of the languages spoken in Mizoram state of India), the primary challenge is the sheer lack of any parallel data. Initial efforts, such as the recent work constructing a large-scale Mizo-English parallel corpus [15], focus on creating these foundational resources from scratch. Similarly, the Indian government’s TDIL program and the Indian Language Corpora Initiative (ILCI) Jha [14] has focused on building a foundational, annotated corpus for key languages. Similarly, Mujadia and Sharma [8] meticulously constructed Hindi-Telugu parallel Corpus using a multi-pronged approach that included human review and back-translation to ensure quality. These foundational efforts, while monumental, are often a first step, as they do not explore whether a smaller, more strategically selected subset of a corpus could achieve comparable or even superior performance. Our work builds on this progress by addressing the next logical challenge: how to most effectively guide such corpus creation by strategically selecting source sentences for cost-efficient model training.
@@ 4.1 Clustering Sentences by LALITA Score
To effectively utilize the LALITA score for data curation, the sentences are grouped into distinct clusters based on their LALITA score (PCA1 values). The Fisher-Jenks algorithm [29] is selected for this clustering task. This statistical method is specifically designed to classify univariate numerical data into contiguous intervals, optimizing for maximum homogeneity within each interval while minimizing variance between intervals. The application of Fisher-Jenks successfully divides the $X_{PCA1}$ [^3] We will be using PCA1 and $X_{PCA1}$ interchangeably values of the sentence feature vectors into four distinct clusters, as shown in Figure 5. The quality of this clustering is assessed using the silhouette score, yielding a value of 0.54, which indicates good separation and cohesion of the clusters. To confirm the robustness of this method, other clustering options, including k-means and Gaussian Mixture Models (GMM), are applied to $X_{PCA2}$ through $X_{PCA10}$ (additional principal components). Nonetheless, these alternative approaches consistently yield substantially lower silhouette scores (below 0.1), reinforcing the superiority of clustering based on the $X_{PCA1}$. Following the clustering of the Filtered English-Hindi Parallel Sentences (FPS), an analysis is conducted on the distribution of source sentences across four clusters, as illustrated in Figure 10. Distinct boundaries are observed between the clusters, with the majority of sentences residing in Cluster 2. A progressive increase in structural complexity among the clusters with LALITA score:
@@
We analyze the cluster-wise distribution of English sentences in the FLORES test set. Figure 8 shows that the sentences in FLORES include a significant number of complex sentences. However, this prompts the question of the nature of the training data, which exhibits a more “uniform distribution”. This offers insight into the disparities between the training and test distributions. It doesn’t imply that a distribution similar to Figure 8 leads to optimum performance, as the system with a comparable distribution in our configuration (10_10_40_40) does not perform as effectively as other configurations (Figure 25). It simply signifies that more consideration is required for the complexity profile of the training data for an effective translation system.
@@ 6.3 Data Filtering
The sheer scale of modern corpora necessitates robust data filtering techniques to remove noise. Shared tasks like Sloto et al. [48] and previous iterations in the WMT have given a platform to filter and curate parallel text. While some efforts rely on manual and heuristic-based filtering, such as in Baturova et al. [49], more automated approaches have emerged. Work like Aulamo et al. [18] provide a wide range of basic filters based on heuristics and metadata. However, the effectiveness of such tools relies on a user’s ability to configure a winning combination of filters, rather than providing a single, principled metric for assessing a sentence’s intrinsic value. Our LALITA framework fills this gap by offering a scientifically grounded, multi-faceted score for linguistic complexity that moves beyond these ad-hoc combinations.
@@ 6.5 Domain Adaptation
The LALITA framework introduces a unique and proactive data curation methodology that focuses exclusively on the linguistic and lexical attributes of the source sentences. This source-centric approach is a critical differentiator, making LALITA particularly applicable in scenarios where target-side quality assessment is difficult or not available. By providing a principled method for curating linguistically complex data, LALITA offers a direct solution to the data quality challenges identified by Khayrallah and Koehn [52]. Our work ensures that NMT models are trained on high-quality data that provides a rich learning signal, thereby improving robustness and performance in a way that simply adding more, uncurated data cannot achieve. Unlike reactive methods that merely clean data or expand its volume, LALITA proactively curates data based on its inherent linguistic value. This approach to data management could lead to new research directions, shifting the focus from the quantity of data to its quality and characteristics.
@@ 7.1 Computational Resource Constraints
The exploratory nature of this research, particularly the exhaustive search for optimal configurations across various dataset sizes and cluster distributions, proved to be computationally intensive and resource-consuming. While the results are compelling, the extensive computational requirements pose a practical barrier to fully exploring the optimal parameter space for LALITA across every possible language pair and resource scenario. This highlights a scalability challenge for comprehensive exploration, suggesting that finding the absolute best configuration might be computationally prohibitive without further optimization of the search process itself.
@@ 8.2 Future Work: Expanding LALITA’s Horizons
  • Enhanced Computational Efficiency: Developing more computationally efficient methods for searching optimal LALITA configurations is a priority. This could involve exploring advanced optimization algorithms or meta-learning approaches to accelerate the discovery of ideal data curation strategies.
@@
  • Optimizing LLM Pretraining with the LALITA Score: Building on this research, a promising future direction is to integrate the LALITA framework with the second-stage training methodology for Large Language Models (LLMs) [59]. Given that LLM pretraining is resource-intensive and “quality trumps over quantity” in data selection, using the LALITA score to proactively curate a high-quality, complex dataset could reduce the vast data volumes and computational costs currently required. By exposing LLMs to a richer variety of structurally complex linguistic constructions, the LALITA score could enhance their ability to learn nuanced grammar and semantics, improving generalization and preventing homogenization to common patterns. Furthermore, it offers a quantifiable metric for strategic data curation and augmentation, allowing for the strategic inclusion of complex synthetic data to address specific linguistic deficiencies in the training corpus.
@@ D.2 Three Clusters
With Skwed Configuration having 70% data drawn from one cluster for training data, we found that systems with cluster 3 having higher sentence percentage (15_15_0_70, 0_15_15_70, 15_0_15_70) performed comparable to FPS, while outperforming other systems. On other side we saw that absence of sentences from Cluster 3 for systems underperformed consistently.
@@ D.3 Two Clusters
  • Skwed sampling (Figure 30): Sampling Majorly from one cluster (75%) while rest (25%) from other cluster.

### v2

@@ Abstract
Data curation is a critical yet under-researched step in the machine translation (MT) training paradigm. To train translation systems, data acquisition relies primarily on human translations and digital parallel sources or, to a limited degree, synthetic generation. But, for low-resource languages, human translation to generate sufficient data is prohibitively expensive. Therefore, it is crucial to develop a framework that screens source sentences to form efficient parallel text, ensuring optimal MT system performance in low-resource environments. We approach this by evaluating English-Hindi bi-text to determine effective sentence selection strategies for optimal MT system training. Our extensively tested framework, (Lexical And Linguistically Informed Text Analysis) LALITA, targets source sentence selection using lexical and linguistic features to curate parallel corpora. We find that by training mostly on complex sentences from both existing and synthetic datasets, our method significantly improves translation quality. We test this by simulating low-resource data availability with curated datasets of 50K to 800K English sentences and report improved performances on all data sizes. LALITA demonstrates remarkable efficiency, reducing data needs by more than half across multiple languages (Hindi, Odia, Nepali, Norwegian Nynorsk, and German). This approach not only reduces MT systems training cost by reducing training data requirement, but also showcases LALITA’s utility in data augmentation.
@@ 1 Introduction
To address this scarcity, foundational efforts have focused on creating and managing large-scale corpora from the ground up. For many low-resource languages, like Mizo (one of the languages spoken in Mizoram state of India), the primary challenge is the sheer lack of any parallel data. Initial efforts, such as the recent work constructing a large-scale Mizo-English parallel corpus [15], focus on creating these foundational resources from scratch. Similarly, the Indian government’s TDIL program and the Indian Language Corpora Initiative (ILCI) Jha [14] have focused on building a foundational, annotated corpus for key languages. Similarly, Mujadia and Sharma [8] meticulously constructed Hindi-Telugu parallel Corpus using a multi-pronged approach that included human review and back-translation to ensure quality. These foundational efforts, while monumental, are often a first step, as they do not explore whether a smaller, more strategically selected subset of a corpus could achieve comparable or even superior performance. Our work builds on this progress by addressing the next logical challenge: how to most effectively guide such corpus creation by strategically selecting source sentences for cost-efficient model training.
@@ 4.1 Clustering Sentences by LALITA Score
To effectively utilize the LALITA score for data curation, the sentences are grouped into distinct clusters based on their LALITA score (PCA1 values). The Fisher-Jenks algorithm [29] is selected for this clustering task. This statistical method is specifically designed to classify univariate numerical data into contiguous intervals, optimising for maximum homogeneity within each interval while minimizing variance between intervals. The application of Fisher-Jenks successfully divides the $X_{PCA1}$ [^3] We will be using PCA1 and $X_{PCA1}$ interchangeably values of the sentence feature vectors into four distinct clusters, as shown in Figure 5. The quality of this clustering is assessed using the silhouette score, yielding a value of 0.54, which indicates good separation and cohesion of the clusters. To confirm the robustness of this method, other clustering options, including k-means and Gaussian Mixture Models (GMM), are applied to $X_{PCA2}$ through $X_{PCA10}$ (additional principal components). Nonetheless, these alternative approaches consistently yield substantially lower silhouette scores (below 0.1), reinforcing the superiority of clustering based on the $X_{PCA1}$. Following the clustering of the Filtered English-Hindi Parallel Sentences (FPS), an analysis is conducted on the distribution of source sentences across four clusters, as illustrated in Figure 10. Distinct boundaries are observed between the clusters, with the majority of sentences residing in Cluster 2. A progressive increase in structural complexity among the clusters with LALITA score:
@@
We analyse the cluster-wise distribution of English sentences in the FLORES test set. Figure 8 shows that the sentences in FLORES include a significant number of complex sentences. However, this prompts the question of the nature of the training data, which exhibits a more “uniform distribution”. This offers insight into the disparities between the training and test distributions. It doesn’t imply that a distribution similar to Figure 8 leads to optimum performance, as the system with a comparable distribution in our configuration (10_10_40_40) does not perform as effectively as other configurations (Figure 25). It simply signifies that more consideration is required for the complexity profile of the training data for an effective translation system.
@@ 6.3 Data Filtering
The sheer scale of modern corpora necessitates robust data filtering techniques to remove noise. Shared tasks like Sloto et al. [48] and previous iterations in the WMT have given a platform to filter and curate parallel text. While some efforts rely on manual and heuristic-based filtering, such as in Baturova et al. [49], more automated approaches have emerged. Work like Aulamo et al. [18] provides a wide range of basic filters based on heuristics and metadata. However, the effectiveness of such tools relies on a user’s ability to configure a winning combination of filters, rather than providing a single, principled metric for assessing a sentence’s intrinsic value. Our LALITA framework fills this gap by offering a scientifically grounded, multi-faceted score for linguistic complexity that moves beyond these ad-hoc combinations.
@@ 6.5 Domain Adaptation
The LALITA framework introduces a unique and proactive data curation methodology that focuses exclusively on the linguistic and lexical attributes of the source sentences. This source-centric approach is a critical differentiator, making LALITA particularly applicable in scenarios where target-side quality assessment is difficult or not available. By providing a principled method for curating linguistically complex data, LALITA offers a direct solution to the data quality challenges identified by Khayrallah and Koehn [52]. Our work ensures that NMT models are trained on high-quality data that provide a rich learning signal, thereby improving robustness and performance in a way that simply adding more, uncurated data cannot achieve. Unlike reactive methods that merely clean data or expand its volume, LALITA proactively curates data based on its inherent linguistic value. This approach to data management could lead to new research directions, shifting the focus from the quantity of data to its quality and characteristics.
@@ 7.1 Computational Resource Constraints
The exploratory nature of this research, particularly the exhaustive search for optimal configurations across various dataset sizes and cluster distributions, proved to be computationally intensive and resource-consuming. While the results are compelling, the extensive computational requirements pose a practical barrier to fully exploring the optimal parameter space for LALITA across every possible language pair and resource scenario. This highlights a scalability challenge for comprehensive exploration, suggesting that finding the absolute best configuration might be computationally prohibitive without further optimisation of the search process itself.
@@ 8.2 Future Work: Expanding LALITA’s Horizons
  • Enhanced Computational Efficiency: Developing more computationally efficient methods for searching optimal LALITA configurations is a priority. This could involve exploring advanced optimisation algorithms or meta-learning approaches to accelerate the discovery of ideal data curation strategies.
@@
  • Optimising LLM Pretraining with the LALITA Score: Building on this research, a promising future direction is to integrate the LALITA framework with the second-stage training methodology for Large Language Models (LLMs) [59]. Given that LLM pretraining is resource-intensive and “quality trumps over quantity” in data selection, using the LALITA score to proactively curate a high-quality, complex dataset could reduce the vast data volumes and computational costs currently required. By exposing LLMs to a richer variety of structurally complex linguistic constructions, the LALITA score could enhance their ability to learn nuanced grammar and semantics, improving generalization and preventing homogenization to common patterns. Furthermore, it offers a quantifiable metric for strategic data curation and augmentation, allowing for the strategic inclusion of complex synthetic data to address specific linguistic deficiencies in the training corpus.
@@ D.2 Three Clusters
With skewed Configuration having 70% data drawn from one cluster for training data, we found that systems with cluster 3 having higher sentence percentage (15_15_0_70, 0_15_15_70, 15_0_15_70) performed comparable to FPS, while outperforming other systems. On other side we saw that absence of sentences from Cluster 3 for systems underperformed consistently.
@@ D.3 Two Clusters
  • skewed sampling (Figure 30): Sampling Majorly from one cluster (75%) while rest (25%) from other cluster.